# Event Ellipsometer: Event-based Mueller-Matrix Video Imaging

### v1

@@ 5 Reconstruction
Figure 4: Synthetic data evaluation result. (a) The rendered images include two materials: blue silicone and brass. (b) The plot shows the error (mean absolute error) of the reconstructed Mueller matrix over the number of iterations. (c) Ground truth Mueller matrix. (d)&(e) Top: the reconstructed Mueller-matrix images for the SVD initialization and the full-stage, respectively. Bottom: a plot of the differentiation of log intensity and fitted line with our method.
@@ 5.2 Spatio-temporal Propagation and Refinement
Figure 5: Assessment of reconstructed Mueller matrix on real data. (a) Evaluation with known optical elements. (b) Measurement on an in-the-wild metal plate induces strong diagonal components.
@@ 7 Results
Figure 6: Photoelasticity analysis. (a) Experimental setup for measuring a gelatine disk in transmission mode. We gradually apply force for the gelatine disk to observe changes in photoelastic properties. (b) Reconstructed Mueller matrix images with different forces, revealing complex stress-dependent polarimetric patterns.
@@
Ellipsometers, as a comprehensive polarization imaging technique for human capture, can reveal hidden properties of the human face and hair that are difficult to analyze in conventional imaging and non-ellipsometric polarization imaging [19, 2]. In Figure 8, we demonstrate the Mueller-matrix reconstruction of dynamic facial expressions and hair movements, revealing intricate polarization properties. The reconstructed Mueller matrix for the face exhibits strong diagonal components associated with specular reflection, while other regions display weakly polarized reflection with a dependency on surface normal. The hair result shows polarization property in the specular highlight regions.
@@ 8 Discussion
The ideal setting of an event camera is to have a low contrast threshold and a short refractory period, to capture subtle intensity changes. However, this produces many events, exceeding the transmission bandwidth, resulting in lost or delayed events [21]. In contrast, when the contrast threshold is high, weakly polarized phenomena cannot be detected as events. Given this trade-off, we defined a region of interest in the event sensor and empirically configured the camera settings to find the best parameters.

### v2

@@ 5 Reconstruction
Figure 4: Synthetic data evaluation result. (a) The rendered images include two materials: blue silicone and brass. (b) The plot shows the error (mean absolute error) of the reconstructed Mueller matrix over the number of iterations. (c) Ground truth Mueller matrix. (d)&(e) Top: the reconstructed Mueller-matrix images for the SVD initialization and the full-stage, respectively. Pixels with insufficient event counts are visualized in white. Bottom: a plot of the differentiation of log intensity and fitted line with our method.
@@ 5.2 Spatio-temporal Propagation and Refinement
Figure 5: Assessment of reconstructed Mueller matrix on real data. (a) Evaluation with known optical elements. We show the corresponding mean squared errors (MSEs). (b) Measurement on an in-the-wild metal plate induces strong diagonal components.
@@ 7 Results
Figure 6: Photoelasticity analysis. (a) Experimental setup for measuring a gelatine disk in transmission mode. We gradually apply force for the gelatine disk to observe changes in photoelastic properties. (b) Reconstructed Mueller matrix images with different forces, revealing complex stress-dependent polarimetric patterns. Notably, as force increases, the Mueller matrix image shows denser fringe patterns.
@@
Ellipsometers, as a comprehensive polarization imaging technique for human capture, can reveal hidden properties of the human face and hair that are difficult to analyze in conventional imaging and non-ellipsometric polarization imaging [19, 2]. In Figure 8, we demonstrate the Mueller-matrix reconstruction of dynamic facial expressions and hair movements, revealing intricate polarization properties. The reconstructed Mueller matrix for the face exhibits strong diagonal components associated with specular reflection, while the non-diagonal regions display weakly polarized reflection with a dependency on surface normal along the face edges. The hair result shows polarization property in the specular highlight regions.
@@ 8 Discussion
The ideal setting of an event camera is to have a low contrast threshold and a short refractory period, to capture subtle intensity changes. However, this produces many events, exceeding the transmission bandwidth, resulting in lost or delayed events [21]. In contrast, when the contrast threshold is high, weakly polarized phenomena cannot be detected as events. Given this trade-off, we defined a region of interest (ROI) in the event sensor and empirically configured the camera settings to find the best parameters.
@@ 9 Conclusion
Acknowledgments
Seung-Hwan Baek was supported by Korea NRF (RS-2023-00211658, RS-2024-00438532, 2022R1A6A1A03052954, RS-2024-00437866), and Samsung Research Funding & Incubation Center of Samsung Electronics under Project Number SRFC-IT1801-52.